OpenStudy (anonymous):

[1-sin(theta)] [1+sin(theta)]=cos^2(theta)

Nnesha (nnesha):

ello o^_^o foil!!!

OpenStudy (anonymous):

Hey! Okay

OpenStudy (anonymous):

I+sine-sine-sine^2=cos^e

Nnesha (nnesha):

 bec there is negative so you can take square of first term and square of last term for example \[\huge\rm x^2 -y^2 =(x-y)(x+y)\]